OpenStudy (anonymous):

Find two solutions of each equation. Express the results in radians.? 1.) CosX= (square root of 2) / 2 2.) CosX= -(square root of 2) / 2 3.) TanX= 1 4.) CotX= -(square root of 3)

OpenStudy (anonymous):

1) cos(x)=√2/2 arccos(√2/2)=x x= π/4 or 7π/4 2) cos(x)=-√2/2 arccos(-√2/2)=x x= 3π/4 or 5π/4 3) tan(x)=1 arctan(1)=x x= π/4, 5π/4 4) cot(x)=-√3 1/tan(x)=-√3 -√3*tan(x)=1 tan(x)=-√3/3 x= 5π/6 and 11π/6

OpenStudy (anonymous):

awesome thanks !